# captionUnknown document class (or package)

### v1

@@ 4.3.2 Training an LLM for multiple 3D tasks
LLMs can be fine-tuned to incorporate multiple 3D tasks using different strategies discussed in Sec. 2.2.3. LEO [270], and 3DMIT [269] use low-rank adaptation (LoRA) for fine-tuning. As a result, the total trainable parameters, including those of alignment modules and 3D encoders, amount to less than 10% of the original LLMs’ parameters, significantly enhancing training efficiency. Chat-3D [171], LL3DA [19], Chat-3D v2 [172], LiDAR-LLM [271], and MultiPLY [24] adopt adaptive fine-tuning. Specifically, these models include modules that align spatial information in 3D scenes with language, e.g. a transformer layer, to capture object relationships. These modules, along with the pre-trained 3D encoders and LLMs, are fine-tuned for alignment. 3D-LLM [153], Scene-LLM [266], Point-LLM [267], and GPT4Point [268] employ layer freezing. By freezing most LLM layers and fine-tuning certain layers like the embedding layer, this strategy preserves linguistic capabilities while improving 3D understanding. Lastly, Agent3D-Zero [17] uses prompt tuning, a training-free method for guiding LLMs toward 3D task comprehension. This approach utilizes tailored prompts, adding grid lines and tick marks onto a BEV image of 3D scene, which help the 2D VLMs understand the 3D geometry.
@@
In contrast to these single-stage alignment methods, LiDAR-LLM [271], Chat-3D [171], and Chat-3D v2 [172] each employ a 2-stage 3D-language alignment process. LiDAR-LLM [271] focuses on enhancing both local and global scene perception through 3D captioning tasks in two phases: first, concentrating on single-view captions and then expanding to panoramic scene descriptions. They develop instance-level perception capabilities through a blend of captioning and grounding tasks. Chat-3D [171] first aligns 3D objects with text using datasets for 3D object classification [296, 293, 297] aiming to maximize the cosine similarity between mapped 3D object features and object category word embeddings by updating only the alignment module. In the second stage of scene-level alignment, it utilizes ScanRefer [218] to enable captioning capabilities, updating an additional transformer layer specifically to model spatial relationships of objects. Similarly, Chat-3D v2 [172] incorporates object-level and scene-level alignments, with the second stage additionally training a position embedding layer. For training efficiency, both LL3DA [19] and 3DMIT [269] skip the alignment stages and focus solely on the stage of instruction tuning described below.
Nearly all multitask learning methods ultimately require the ability to complete various 3D tasks based on instructions. Thus, as the final stage of training, each method typically uses their own constructed multitask instruction-following dataset for instruction fine-tuning [^1] Instruction fine-tuning refers to the process of further fine-tuning LLMs on a dataset consisting of (INSTRUCTION, OUTPUT) pairs [298]. Since all task outputs are unified into textual form, the training loss used is the standard autoregressive loss used in LLMs. This stage usually involves jointly training the alignment module and the LLM. An exception is Agent3D-Zero [17], which completes various 3D tasks by feeding GPT4V with 2D images from different viewpoints, thus it does not require any training.
@@ 4.4 LLMs as 3D Multi-Modal Interfaces
4.5 LLMs for Embodied Agents
@@
LEO [270] and LLM-Planner [12] utilize LLMs to generate step-by-step plans which they adjust dynamically based on environmental perceptions. LEO [270] emphasizes scene-aware planning grounded in the current scene’s configurations, whereas LLM-Planner [12] adopts GPT3 [127] to divide planning into high-level subgoals and low-level actions, and replan when the agent gets stuck during task execution. 3D-VLA [276] integrates 3D perception, reasoning, and action through a generative world model. It focuses on enhancing planning capabilities by utilizing its generative model to predict future state representations (e.g. goal images and point clouds). Agent3D-Zero [17] introduces Set-of-Line Prompting (SoLP) that enhances the VLM’s comprehension of the scene’s geometric aspects by generating diverse observational viewpoints. Specifically, SoLP superimposes grid lines and tick marks onto BEV images, and prompts the VLM to provide more accurate camera positions and orientations, which enables the task of the VLM to understand 3D spatial concepts. UniHSI [277] addresses the task of Human-Scene Interaction (HSI), which involves generating interactions between humans and objects within 3D environments based on input language commands. It uses an LLM as a planner to translate language commands into task plans represented as Chains of Contacts (CoC), a sequence that represents the chronological relationships between human joint points and object positions. While the aforementioned approaches focus on planning within a single scene, SayPlan [275] can handle multiple rooms and floors by (i) leveraging 3D scene graphs for semantic search and (ii) integrating classical path planning with an iterative replanning pipeline for plan refinement.
@@
3D Navigation refers to the ability of an embodied agent to move and orient itself within 3D environments, often based on visual inputs and language instructions. Each of the described methods – LEO [270], Agent3D-Zero [17], LLM-Planner [12], and NaviLLM [11] – implements 3D navigation in different ways. LEO [270] processes egocentric 2D images and object-centric 3D point clouds alongside textual instructions. It generates a sequence of action tokens that correspond to executable navigation commands like ‘move forward’ or ‘turn right’. LEO employs ‘shortest path navigation trials’, which provide a less noisy and more straightforward learning environment compared to human demonstrations. Agent3D-Zero [17] navigates by continuously selecting new viewpoints based on assessments of the environment. It incorporates historical data from previous viewpoints to refine its navigation path towards specific goals, such as finding a printer in an office setting. LLM-Planner [12] employs a hierarchical approach, that first generates high-level plans as sequences of subgoals, which are then translated into a series of primitive actions by a low-level planner. This makes the overall process adaptable to the immediate environment. NaviLLM [11] transforms various embodied navigation tasks into generation problems using schema-based instructions. These instructions include $4$ elements: a task defined by a word sequence, observations of all reachable viewpoints, the history of past visual observations, and output hints that guide action generation (e.g. selecting a direction or object).
@@ 4.6.1 Object-level Generation
In contrast, MeshGPT [282] and PolyGen [280] do not condition the generation on text, but they still adopt an autoregressive approach akin to sequence modelling in LLMs. MeshGPT uses graph convolutions to encode mesh geometry/topology into rich embeddings compressed via Residual Vector Quantization and fed into a GPT-style transformer to autoregressively predict tokens/embeddings for generating meshes with desired properties. PolyGen [280] is an autoregressive transformer-based model of 3D meshes, which utilizes pointer networks. It consists of a vertex model that unconditionally models mesh vertices, and a face model for mesh faces conditioned on input vertices using autoregressive networks to output face indices and vertex coordinates for generating diverse high-quality meshes.
@@ 4.6.2 Scene-scale Generation
GALA-3D [25] first generates coarse layouts from text using an LLM, then translates them into a 3D Gaussian representation. This representation serves as the foundation for creating detailed 3D content using instance-level text-to-image diffusion priors. It employs compositional optimization to fine-tune the parameters of the layout-guided Gaussians, ensuring the final scene aligns with the text for object placement, scale, and interaction.
Both leverage complementary strengths of LLMs to extract high-level semantic layouts and generative models/optimization to convert these into geometrically and physically plausible 3D scenes.
@@ 4.6.3 Procedural Generation and Manipulation
While Section 4 discussed methods integrating LLMs in 3D tasks, a vast body of research has explored various aspects of 3D understanding through the lens of 2D Vision-Language Models (VLMs). VLMs include much richer visual information that can be linked directly to 3D. This section reviews contributions from a series of recent papers covering language-driven open-world understanding, instance-level understanding, unified end-to-end architectures, spatial reasoning, generation, and beyond.
@@ 5.1 Open-Vocabulary 3D Scene Understanding
Open-vocabulary 3D scene understanding aims to recognize and describe scene elements using natural language descriptions instead of predefined category labels. OpenScene [29] takes a zero-shot approach by predicting dense features for 3D scene points that are co-embedded with CLIP’s text and image pixel embeddings in a shared feature space, enabling task-agnostic training and open-vocabulary querying to identify objects, materials, affordances, activities, and room types. CLIP-FO3D [304] follows a similar approach, modifying CLIP to extract dense pixel features from 3D scenes which are projected to point clouds, then training a 3D model via distillation to transfer CLIP’s knowledge. Semantic Abstraction [305] extracts relevancy maps from CLIP as abstracted object representations to generalize to novel semantics, vocabularies, and domains. Open-Fusion [306] combines the SEEM [307] vision-language model with TSDF 3D mapping for real-time open-vocabulary scene creation and querying, utilizing region-based embeddings and confidence maps.
Approaches such as PLA [308] and RegionPLC [309] leverage contrastive learning to combine captions with 2D and 3D data modalities to associate visual and semantic information. PLA [308] uses 3D-caption pairs and contrastive learning to associate multi-view images with captions for learning visual-semantic representations, while RegionPLC [309] proposes region-aware contrastive learning by combining region-level captions from 2D models mapped to 3D points. OVIR-3D [310] fuses 2D region proposals and text-aligned features from an off-the-shelf 2D detector into 3D instances for efficient open-vocab retrieval. CoDA [311] uses 3D geometry priors from annotated base categories and CLIP’s 2D semantic priors in its 3D Novel Object Discovery (3D-NOD) strategy. Its Discovery-driven Cross-Modal Alignment (DCMA) aligns 3D and image/text features for novel object localization and classification.
Instance-level scene understanding works such as OpenMask3D [312] and Open3DIS [313] utilize predicted class-agnostic 3D instance masks and 2D segment-level CLIP embeddings to enable open-vocabulary 3D instance segmentation. OpenIns3D [314] achieves open-vocabulary understanding without aligned images, using a “Mask-Snap-Lookup” pipeline that predicts 3D mask proposals, generates synthetic scene images, and assigns categories to masks via a language module. Rozenberszki et al.[315] propose to leverage CLIP features to ground 3D feature learning for 3D semantic and instance segmentation.
Language grounding with NeRFs has shown promising results in open-vocabulary scene understanding. Several methods, such as DFF [316], LERF [30], VL-Fields [317], and 3D-OVS [318], distill the knowledge of 2D feature extractors like DINO or CLIP into 3D feature fields by minimizing the error of volume-rendered features with respect to the 2D features, enabling query-based local editing and grounding language into neural implicit representations. LERF [30] optimizes a dense, scale-conditioned 3D language field by volume rendering CLIP embeddings. LangSplat [319] and N2F2 [320] demonstrate efficient open-vocabulary querying and interaction within a 3D Gaussian Splatting representation by leveraging hierarchical supervision and multiscale feature fields.
@@ 5.2 Text-Driven 3D Generation
Section 4.6 covered methods that used LLMs for 3D generation. Here, we survey text-to-3D generation methods that leverage guidance from 2D VLMs [174] and text-to-image diffusion models [321, 322] using differentiable rendering. Early works such as DreamFields [323], CLIP-Mesh [32], CLIP-Forge [324] and Text2Mesh [325] explore zero-shot 3D generation guided by CLIP.
DreamFusion [31] introduced Score Distillation Sampling (SDS) where the parameters of a 3D representation are optimized by making its renderings from arbitrary perspectives appear highly realistic, as evaluated by a pre-trained 2D diffusion model. It uses the text-to-image Imagen model [322] to optimize a NeRF representation via SDS. Magic3D [326] proposes a two-stage framework: generating a coarse model with a low-resolution diffusion prior and sparse 3D hash grid, then optimizing a textured 3D mesh model with an efficient differentiable renderer and high-resolution latent diffusion model [321]. Fantasia3D [327] disentangles geometry and appearance, using a hybrid DMTet [328] representation and spatially varying BRDFs. ProlificDreamer [329] introduced Variational Score Distillation (VSD), a particle-based framework treating 3D parameters as random variables to improve fidelity and diversity. Dream3D [330] utilizes explicit 3D shape priors and text-to-image diffusion models for enhanced text-guided 3D synthesis. MVDream [331] employs a multi-view consistent diffusion model trainable on few-shot data for personalized generation. Text2NeRF [332] combines NeRF representations with pre-trained text-to-image diffusion models to generate diverse indoor/outdoor 3D scenes from language. Apart from simultaneously generating geometries and appearance, several studies explore the possibility of solely synthesizing the textures based on given geometries [333, 334, 335].
For human avatars, AvatarCraft [336] uses diffusion models to guide neural implicit field geometry/texture learning from text prompts. Additionally, it enables animating these human avatars by deforming the neural implicit field with an explicit warping field that maps the target human mesh to a template human mesh. AvatarCLIP [337] proposes a zero-shot CLIP-supervised framework for 3D avatar generation, geometry sculpting, texture mapping and motion synthesis from text. CG-HOI [338] uses diffusion models to characterize dynamic human-object interactions from text. GenZI [339] distills information about human interactions via a pre-trained vision-language model, producing zero-shot 3D human-scene interaction synthesis from text prompts.
Exploring compositional generation, CG3D [340] generates scalable 3D scenes using explicit 3D Gaussian radiance fields by compositing individual objects without bounding boxes. Po et al. [341] introduce locally conditioned diffusion for granular scene control via text prompts and bounding boxes. GraphDreamer [342] generates compositional scenes from scene graphs by decomposing them into global-local descriptions to optimize object SDFs.
Overall, these approaches combine diffusion models, vision-language models, neural representations, and 3D priors for text-to-3D generation of objects, avatars, and scenes.

### v2

@@ 4.3.3 Training Strategy for Multi-Task Learning
LLMs can be fine-tuned to incorporate multiple 3D tasks using different strategies discussed in Sec. 2.2.3. LEO [152], MiniGPT-3D [159], and 3DMIT [157] use low-rank adaptation (LoRA) for fine-tuning. As a result, the total trainable parameters, including those of alignment modules and 3D encoders, amount to less than 10% of the original LLMs’ parameters, significantly enhancing training efficiency. Chat-3D [80], LL3DA [19], Chat-3D v2 [81], LiDAR-LLM [155], and MultiPLY [24] adopt adaptive fine-tuning. Specifically, these models include modules that align spatial information in 3D scenes with language, e.g. a transformer layer, to capture object relationships. These modules, along with the pre-trained 3D encoders and LLMs, are fine-tuned for alignment. 3D-LLM [115], Scene-LLM [147], Point-LLM [153], and GPT4Point [154] employ layer freezing. By freezing most LLM layers and fine-tuning certain layers like the embedding layer, this strategy preserves linguistic capabilities while improving 3D understanding. Lastly, Agent3D-Zero [17] uses prompt tuning, a training-free method for guiding LLMs toward 3D task comprehension. This approach utilizes tailored prompts, adding grid lines and tick marks onto a BEV image of 3D scene, which help the 2D VLMs understand the 3D geometry.
@@
By contrast, other works [80, 81, 155, 159] employ a multi-stage alignment process, with each stage designed to enhance different aspects of the LLM’s 3D capabilities. LiDAR-LLM [155] focuses on enhancing both local and global scene perception through 3D captioning tasks in two phases: first, concentrating on single-view captions and then expanding to panoramic scene descriptions. They develop instance-level perception capabilities through a blend of captioning and grounding tasks. Chat-3D [80] first aligns 3D objects with text using datasets for 3D object classification [204, 201, 205], aiming to maximize the cosine similarity between mapped 3D object features and object category word embeddings. In the second stage, it utilizes ScanRefer [101] to enable captioning capabilities, updating an additional transformer layer specifically to model spatial relationships of objects. Similarly, Chat-3D v2 [81] incorporates object-level and scene-level alignments, with the second stage training an additional position embedding layer. MiniGPT-3D [159] additionally fine-tunes a ‘Mixture of Query’ module to enhance their model’s 3D perception ability from multiple perspectives. GreenPLM [161] propose to only align text caption of 3D data to LLM at the first two stages and then achieve object-level alignment by tuning the MLP layer. For training efficiency, both LL3DA [19] and 3DMIT [157] skip the alignment stages and focus solely on the stage of instruction tuning described below.
Nearly all multi-task learning methods ultimately require the ability to complete various 3D tasks based on instructions. Thus, as the final stage of training, each method typically uses their own constructed multitask instruction-following dataset for instruction fine-tuning [^1] Instruction fine-tuning refers to the process of further fine-tuning LLMs on a dataset consisting of (INSTRUCTION, OUTPUT) pairs [206]. Since all task outputs are unified into textual form, the training loss used is the standard auto-regressive loss used in LLMs, except that Ross3D [170] uses an additional reconstructive loss. This stage usually involves jointly training the alignment module and the LLM. An exception is Agent3D-Zero [17], which completes various 3D tasks by feeding GPT-4V with 2D images from different viewpoints, thus it does not require any training.
@@ 4.4 LLMs as 3D Multi-Modal Interfaces
4.5 LLMs for 3D Embodied Agents
@@
LEO [152] and LLM-Planner [12] utilize LLMs to generate step-by-step plans which they adjust dynamically based on environmental perceptions. LEO [152] emphasizes scene-aware planning grounded in the current scene’s configurations, whereas LLM-Planner [12] adopts GPT-3 [70] to divide planning into high-level subgoals and low-level actions, and replan when the agent gets stuck during task execution. 3D-VLA [182] integrates 3D perception, reasoning, and action through a generative world model. It focuses on enhancing planning capabilities by utilizing its generative model to predict future state representations (e.g. goal images and point clouds). Agent3D-Zero [17] introduces Set-of-Line Prompting (SoLP) that enhances the VLM’s comprehension of the scene’s geometric aspects by generating diverse observational viewpoints. Specifically, SoLP superimposes grid lines and tick marks onto BEV images, and prompts the VLM to provide more accurate camera positions and orientations, which enables the task of the VLM to understand 3D spatial concepts. UniHSI [181] addresses the task of Human-Scene Interaction (HSI), which involves generating interactions between humans and objects within 3D environments based on input language commands. It uses an LLM as a planner to translate language commands into task plans represented as Chains of Contacts (CoC), a sequence that represents the chronological relationships between human joint points and object positions. While the aforementioned approaches focus on planning within a single scene, SayPlan [180] can handle multiple rooms and floors by (i) leveraging 3D scene graphs for semantic search and (ii) integrating classical path planning with an iterative replanning pipeline for plan refinement. NLMap [214] proposes an open-vocabulary, queryable scene map for LLM planners, enabling them to see and query available objects before generating plans.
@@
3D Navigation refers to the ability of an embodied agent to move and orient itself within 3D environments, often based on visual inputs and language instructions. Each of the described methods – LEO [152], Agent3D-Zero [17], LLM-Planner [12], and NaviLLM [11] – implements 3D navigation in different ways. LEO [152] processes egocentric 2D images and object-centric 3D point clouds alongside textual instructions. It generates a sequence of action tokens that correspond to executable navigation commands like ‘move forward’ or ‘turn right’. LEO employs ‘shortest path navigation trials’, which provide a less noisy and more straightforward learning environment compared to human demonstrations. Agent3D-Zero [17] navigates by continuously selecting new viewpoints based on assessments of the environment. It incorporates historical data from previous viewpoints to refine its navigation path towards specific goals, such as finding a printer in an office setting. LLM-Planner [12] employs a hierarchical approach that first generates high-level plans as sequences of subgoals, which are then translated into a series of primitive actions by a low-level planner. This makes the overall process adaptable to the immediate environment. NaviLLM [11] transforms various embodied navigation tasks into generation problems using schema-based instructions. These instructions include $4$ elements: a task defined by a word sequence, observations of all reachable viewpoints, the history of past visual observations, and output hints that guide action generation (e.g. selecting a direction or object).
@@ 4.6.1 Object-level Generation
In contrast, MeshGPT [185], LLaMA-Mesh [188], and PolyGen [183] do not condition the generation on text, but they still adopt an autoregressive approach akin to sequence modelling in LLMs. MeshGPT uses graph convolutions to encode mesh geometry/topology into rich embeddings compressed via Residual Vector Quantization and fed into a GPT-style transformer to autoregressively predict tokens/embeddings for generating meshes with desired properties. PolyGen [183] is an autoregressive transformer-based model of 3D meshes, which utilizes pointer networks. It consists of a vertex model that unconditionally models mesh vertices, and a face model for mesh faces conditioned on input vertices using autoregressive networks to output face indices and vertex coordinates for generating diverse high-quality meshes. LLaMA-Mesh [188] represents 3D meshes in a purely textual form by expressing vertex coordinates and face indices as plain text sequences, with coordinate values discretized into integers for compactness, allowing the direct use of pretrained LLM tokenizers and architectures without introducing a mesh-specific vocabulary.
@@ 4.6.2 Scene-scale Generation
GALA-3D [25] first generates coarse layouts from text using an LLM, then translates them into a 3D Gaussian representation. This representation serves as the foundation for creating detailed 3D content using instance-level text-to-image diffusion priors. It employs compositional optimization to fine-tune the parameters of the layout-guided Gaussians, ensuring the final scene aligns with the text for object placement, scale, and interaction. Both leverage complementary strengths of LLMs to extract high-level semantic layouts and generative models/optimization to convert these into geometrically and physically plausible 3D scenes.
@@ 4.6.3 Procedural Generation and Manipulation
While Section 4 discussed methods integrating LLMs in 3D tasks, a vast body of research has explored various aspects of 3D understanding through the lens of 2D Vision-Language Models (VLMs). VLMs include richer visual priors that can be linked to 3D representations. This section reviews contributions from recent papers spanning open-world understanding, instance-level recognition, unified end-to-end architectures, spatial reasoning, and generation.
@@ 5.1 Open-Vocabulary 3D Scene Understanding
Open-vocabulary 3D scene understanding recognizes and describes scene elements using natural language instead of predefined category labels. OpenScene [217] adopts a zero-shot strategy by predicting dense 3D features co-embedded with CLIP’s text and image embeddings in a shared space, enabling task-agnostic querying of objects, materials, affordances, and room types. CLIP-FO3D [218] similarly distills dense pixel features from CLIP to 3D point clouds via feature projection and training a 3D model to absorb CLIP knowledge. Semantic Abstraction [219] extracts CLIP relevance maps as abstract object representations to generalize across vocabularies and domains. Open-Fusion [220] integrates SEEM [221] with TSDF 3D mapping for real-time open-vocabulary scene creation using region-based embeddings and confidence maps.
PLA [222] and RegionPLC [223] use contrastive learning to associate captions with 2D and 3D modalities. PLA aligns multi-view images with captions via 3D-caption contrastive pairs, while RegionPLC aligns region-level 2D captions to 3D points. OVIR-3D [224] fuses 2D region proposals with CLIP-aligned features to support efficient open-vocabulary 3D instance retrieval. CoDA [225] incorporates annotated 3D geometry priors with 2D semantics from CLIP to discover novel objects using Discovery-driven Cross-Modal Alignment (DCMA).
For instance-level understanding, OpenMask3D [197] and Open3DIS [226] use class-agnostic 3D instance masks and 2D segment-level CLIP embeddings for open-vocabulary 3D instance segmentation. OpenIns3D [227] bypasses image alignment by predicting 3D mask proposals, rendering synthetic images, and assigning mask labels via a language model. Rozenberszki et al. [228] align 3D features with CLIP for semantic and instance-level segmentation. Diff2Scene [229] distills semantic knowledge from frozen text-to-image diffusion models, using geometry- and saliency-aware masks to enable label-free open-vocabulary 3D segmentation and grounding.
Language grounding with NeRFs has shown promise in open-vocabulary 3D understanding. Methods like DFF [230], LERF [231], VL-Fields [232], and 3D-OVS [233] distill 2D VLM knowledge into implicit 3D feature fields by minimizing rendering error w.r.t. 2D features. LERF further learns a scale-conditioned language field over 3D space. LangSplat [234] and N2F2 [235] extend this to 3D Gaussian Splatting, enabling efficient querying via hierarchical supervision and multiscale feature fields.
@@ 5.2 Text-Driven 3D Generation
Section 4.6 reviewed LLM-guided generation; here, we survey methods leveraging 2D VLMs [82] and text-to-image diffusion models [236, 237] with differentiable rendering. Early methods like DreamFields [238], CLIP-Mesh [239], CLIP-Forge [240], and Text2Mesh [241] explore CLIP-guided zero-shot 3D optimization.
DreamFusion [92] introduced Score Distillation Sampling (SDS), optimizing NeRF parameters so rendered views are realistic under a text-to-image diffusion model (Imagen [237]). Magic3D [242] adds a two-stage pipeline - coarse NeRF generation followed by mesh optimization using high-res latent diffusion. Fantasia3D [243] disentangles geometry and appearance via a DMTet [244] hybrid with spatially varying BRDFs. ProlificDreamer [245] introduces Variational SDS (VSD), treating 3D parameters as particles for more diverse generation. Dream3D [246] enhances synthesis using explicit 3D shape priors. MVDream [247] trains multi-view consistent diffusion on few-shot data. Text2NeRF [248] uses pre-trained diffusion models to guide NeRF generation from language. CrossOver [249] learns a modality-agnostic space across images, point clouds, CAD models, floorplans, and text, enabling cross-modal scene retrieval and generation.
Some works target text-guided texture synthesis only [250, 251, 252]. For avatars, AvatarCraft [253] uses diffusion to learn geometry/texture from text and animate avatars via warping fields. AvatarCLIP [254] provides a zero-shot CLIP-supervised framework for geometry, texture, and motion synthesis. GenZI [255] and CG-HOI [256] model human-object interactions from text.
Compositional approaches like CG3D [257] assemble Gaussian radiance fields from object parts. Po et al. [258] enable local scene control via text and bounding boxes. GraphDreamer [259] composes scenes from graph-derived global-local descriptions using object SDF optimization.